OpenStudy (anonymous):

Mrs. Grijalva hired a landscaping company to landscape her back yard. The company has a standing policy to plant 3 bushes for every tree planted. The company charges $75 for each tree planted and $20 for each bush. Mrs. Grijalva was billed for $810 after the planting was complete. How many bushes did the company plant in Mrs. Grijalva’s back yard? A. 2 B. 6 C. 18 D. 54

OpenStudy (anonymous):

B=3T the bushes cost $20 each, and the trees costs $75 each =$20*B+$75*T the total cost was $810 810=20B+75T 810=20(3T)+75T 810=135T T=6 plug into formula B=3T=3(6)=18

OpenStudy (anonymous):

thxs